OpenStudy (anonymous):

Looking for python help: How do i make a while loop reset and allow the user to input new information

OpenStudy (e.mccormick):

You mean reset the values inside a loop? Probably need some conditional statement (if) to see when this needs to happen.

OpenStudy (anonymous):

ooh ok so you can put an if statement inside of a while ?

OpenStudy (anonymous):

``` print('1. Argentine Peso') print('2. Brazilian Real') print('3. British Pound') print('4. European Euro') print('5. Indian Rupee') print('6. Japanese Yen') print('7. Russian Ruble') print('8. Other \n') usd = 0 usd = eval(input('Enter amount to convert (<= 0 to exit): ')) peso = usd * 8.48 real = usd * 2.46 pound = usd * 0.62 euro = usd * 0.79 rupee = usd * 61.59 yen = usd * 109.49 ruble = usd * 39.45 num = eval(input('Enter Selection: ')) while num == 1: num = num + 1 print('' + str(usd) + ' US Dollars converts to: ' + str(round(peso,2)) + ' Argentine Pesos.') while num == 2: num = num + 1 print('' + str(usd) + ' US Dollars converts to: ' + str(round(real,2)) + ' Brazilian Real.') while num == 3: num = num + 1 print('' + str(usd) + ' US Dollars converts to: ' + str(round(pound,2)) + ' British Pound.') while num == 4: num = num + 1 print('' + str(usd) + ' US Dollars converts to: ' + str(round(euro,2)) + ' European Euro.') while num == 5: num = num + 1 print('' + str(usd) + ' US Dollars converts to: ' + str(round(rupee,2)) + ' Indian Rupee.') while num == 6: num = num + 1 print('' + str(usd) + ' US Dollars converts to: ' + str(round(yen,2)) + ' Japanese Yen.') while num == 7: num = num + 1 print('' + str(usd) + ' US Dollars converts to: ' + str(ruble) + ' Russian Ruble.') while num == 8: num = num + 1 currency = input('Enter current to convert to: ') amount = eval(input('Enter the conversion factor for ' + currency + ': ')) random = usd * amount print('' + str(usd) + ' US Dollars converts to: ' + str(random) + ' ' + currency + '.') ```

OpenStudy (anonymous):

this is currently my code, it runs fine but i need it to only use one of the options at a time, then to restart and allow the user to start over and enter a new number and choose a new option

OpenStudy (anonymous):

@e.mccormick

OpenStudy (e.mccormick):

You can even put a while inside a while.

OpenStudy (anonymous):

hmm arlight, let me give it a shot

OpenStudy (e.mccormick):

Yah, typically what you are doing would be done with ifs or in some languages with a select/case.

OpenStudy (anonymous):

yeah i just figured i was supposed to use while because thats what we were going over in class and thought thats what this assignment was supposed to teach us.

OpenStudy (e.mccormick):

Well, if the entire program is meant to run multiple times, you could use a while loop around it all and then put the if stuff inside it.

OpenStudy (anonymous):

oooh! ok so maybe thats what im supposed to do

OpenStudy (anonymous):

that makes sense!

OpenStudy (anonymous):

thank you, ill use the info you gave me. i gotta go, thanks again

OpenStudy (e.mccormick):

Kk. It is a pretty simple concept, so you should be able to figure it out, but sometimes a diagram helps. http://www.cs.kent.edu/~batcher/CS10051/loops.jpg There is one I found with an image search. The set of algorithmic operations to be performed is the conversion and output stuff.